PLEASE ID APPRECIATE YOU ANSWERING THIS SO SOON...

It’s

add or subtract the following vectors and write down the vector result.

a) 200N-> + 100N-> =

b) 300m -> + <-100m=

c) 25N-> + 7N-> + <- 9N + 12N-> =

d) 6 miles north and 13 miles south =

e) 83m/s east and 51m/s south =



another questions

1) a man walks 2.6 miles east then turns and walks 6.7 miles north. Where is he in relation to his starting point?

2) Find the displacement done by an object towards east if you know that it made 25m displacement towards north and the total result of this situation is 40m.


1
Expert's answer
2021-01-21T16:14:22-0500

a)


"200+100=300\ N"

b)


"300-100=200\ m"

c)


"25+7-9+12=35\ N"

d)


"13-6=7\ south"

e)


"v=\sqrt{83^2+51^2}=97.4\frac{m}{s}"

1)


"s=\sqrt{2.6^2+6.7^2}=7.2\ m"

2)


"d=\sqrt{40^2-25^2}=31.2\ m"